# Orak: A Foundational Benchmark for Training and Evaluating LLM Agents on Diverse Video Games

### v1

@@ I.5 Result for Minecraft
Llama-3.2-1B 0.0$\pm$0.0 9.5
Llama-3.2-3B 0.0$\pm$0.0 9.5
Qwen-2.5-3B 0.0$\pm$0.0 9.5
Qwen-2.5-7B 0.0$\pm$0.0 9.5
Minitron-4B 0.0$\pm$0.0 9.5
Minitron-8B 0.0$\pm$0.0 9.5
GPT-4o-mini 46.0$\pm$7.0 5
GPT-4o 71.0$\pm$7.0 4
o3-mini 75.0$\pm$0.0 2
Gemini-2.5-pro 75.0$\pm$0.0 2
Claude-3.7 75.0$\pm$0.0 2
Deepseek-R1 41.7$\pm$0.0 6
@@
Llama-3B Zero-shot 0.0$\pm$0.0 6
Reflection 0.0$\pm$0.0 6
Planning 0.0$\pm$0.0 6
Ref-Plan 0.0$\pm$0.0 6
GPT-4o Zero-shot 0.0$\pm$0.0 6
Reflection 50.0$\pm$0.0 1.5
Planning 13.0$\pm$0.0 3
Ref-Plan 50.0$\pm$0.0 1.5
@@
Environment. Stardew Valley [61] is a life simulation and farming role-playing game. The player can engage in a variety of daily activities such as farming, fishing, mining, foraging, and socializing with villagers.
@@
We run the game on the Steam platform and use the modding tool SMAPI [84] to extract in-game states and implement custom actions. To send keyboard and mouse inputs, we use the pyautogui library on macOS and the AutoHotkey (AHK) library on Windows, following prior work [31]. Since many in-game actions such as planting or watering crops consist of multiple low-level actions (e.g., move up/down/right/left, switch tool, use tool), we define a set of high-level actions to abstract these into semantically meaningful units. Each high-level action is mapped to a predefined sequence of keyboard inputs and implemented via SMAPI scripts with custom keyboard bindings. We provide the save point used for our experiment to ensure reproducibility.
@@
The objective in the Stardew Valley task is to maximize the amount of money earned during the first 13 in-game days, starting from Day 1. This period is a natural milestone in the game, as the Egg Festival takes place on Spring 13, where players can purchase high-reward crops such as Strawberry seeds. We evaluate performance based on the net profit earned by the end of Spring 13, calculated as the difference between the final gold amount and the initial amount of 200 gold. We normalize the score using a human expert baseline of $1013-200=800$ gold, which we assign a normalized score of 100. Formally, the normalized score is defined as
$\text{Score}=(x_{\text{final}}-x_{\text{start}})/(x_{\text{oracle}}-x_{\text{start}})=(x_{\text{final}}-x_{\text{start}})/813,$
where $x_{\text{final}}$ is the agent’s final gold amount, $x_{\text{start}}=200$ is the starting gold, and $x_{\text{oracle}}=1013$ is the human expert score.
@@
Llama-3.2-1B 0.0$\pm$0.0 9.5
Llama-3.2-3B 0.0$\pm$0.0 9.5
Qwen-2.5-3B 0.0$\pm$0.0 9.5
Qwen-2.5-7B 0.0$\pm$0.0 9.5
Minitron-4B 0.0$\pm$0.0 9.5
Minitron-8B 0.0$\pm$0.0 9.5
GPT-4o-mini 18.9$\pm$32.7 6
GPT-4o 95.7$\pm$5.7 1
o3-mini 64.7$\pm$18.8 4
Gemini-2.5-pro 69.6$\pm$11.9 3
Claude-3.7 77.7$\pm$13.6 2
Deepseek-R1 63.0$\pm$24.6 5
@@
Llama-3B Zero-shot 0.0$\pm$0.0 6.5
Reflection 0.0$\pm$0.0 6.5
Planning 0.0$\pm$0.0 6.5
Ref-Plan 0.0$\pm$0.0 6.5
GPT-4o Zero-shot 40.5$\pm$35.2 3
Reflection 18.3$\pm$5.2 4
Planning 64.6$\pm$23.7 2
Ref-Plan 95.7$\pm$5.7 1
@@
GPT-4o Text 95.7$\pm$5.7 1
Image 0.0$\pm$0.0 8.5
Both 49.2$\pm$26.0 6
Gemini Text 69.6$\pm$11.9 3
Image 8.9$\pm$10.5 7
Both 70.6$\pm$7.1 2
Claude Text 63.0$\pm$24.6 4
Image 0.0$\pm$0.0 8.5
Both 58.5$\pm$1.2 5
@@
Environment. StarCraft II is a real-time strategy game where players gather resources, construct buildings, train units, and command armies to defeat opponents. The environment features a partially observable map, requiring the agent to explore and gather information about the opponent’s actions. For implementation, we adopt the BurnySc2/python-sc2 environment [85], a Python interface widely used in the reinforcement learning (RL) community. Note that the library supports the raw scripted interface without a graphics-based interface. The environment supports various official maps and game modes. For our default evaluation (as in Table 3), we use the ‘Ancient Cistern LE’ map with the agent playing as Protoss against the built-in AI bot (Zerg, Hard difficulty; employing a timing build order strategy). The environment allows testing across different races, maps, and difficulty settings. For instance, the ‘Babylon LE’ is used to assess intra-game generalization in Table 8.
@@
Score=Number of WinsTotal Matches $\text{Score}=\frac{\text{Number of Wins}}{\text{Total Matches Played}}\times 100$
@@
Using these difference-based scores, we then compute Elo ratings for all agents following the Bradley-Terry model [76], following the approach used in StreetFighter III multi-agent evaluation in Section C.3.
@@
Llama-3.2-1B 0.0$\pm$0.0 9.5
Llama-3.2-3B 0.0$\pm$0.0 9.5
Qwen-2.5-3B 0.0$\pm$0.0 9.5
Qwen-2.5-7B 0.0$\pm$0.0 9.5
Minitron-4B 0.0$\pm$0.0 9.5
Minitron-8B 0.0$\pm$0.0 9.5
GPT-4o-mini 75.0$\pm$50.0 3
GPT-4o 100.0$\pm$0.0 1.5
o3-mini 25.0$\pm$50.0 6
Gemini-2.5-pro 100.0$\pm$0.0 1.5
Claude-3.7 50.0$\pm$57.7 4.5
Deepseek-R1 50.0$\pm$57.7 4.5
@@
Llama-3B Zero-shot 0.0$\pm$0.0 6
Reflection 0.0$\pm$0.0 6
Planning 0.0$\pm$0.0 6
Ref-Plan 0.0$\pm$0.0 6
GPT-4o Zero-shot 50.0$\pm$57.7 2.5
Reflection 0.0$\pm$0.0 6
Planning 50.0$\pm$57.7 2.5
Ref-Plan 100.0$\pm$0.0 1
@@
GPT-4o Text 100.0$\pm$0.0 2
Both 50.0$\pm$57.7 5
Gemini Text 100.0$\pm$0.0 2
Both 100.0$\pm$0.0 2
Claude Text 50.0$\pm$57.7 5
Both 50.0$\pm$57.7 5
@@
The game runs on the Steam platform, and we use modding tools to enable communication between the game and an external agent. In particular, we use BaseMod [86] and ModTheSpire [87], along with a modified version of CommunicationMod [88], which enables state extraction and action input via standard input/output streams. Our modified version also extracts detailed in-game information, including full card and relic descriptions.
@@
[…]
@@
Llama-3.2-1B 0.0$\pm$0.0 10
Llama-3.2-3B 0.0$\pm$0.0 10
Qwen-2.5-3B 0.0$\pm$0.0 10
Qwen-2.5-7B 5.0$\pm$0.0 6
Minitron-4B 0.0$\pm$0.0 10
Minitron-8B 0.0$\pm$0.0 10
GPT-4o-mini 3.3$\pm$2.9 7
GPT-4o 23.6$\pm$22.1 3
o3-mini 15.0$\pm$0.0 4.5
Gemini-2.5-pro 51.9$\pm$31.9 1
Claude-3.7 15.0$\pm$0.0 4.5
Deepseek-R1 24.9$\pm$17.1 2
@@
Llama-3B Zero-shot 0.0$\pm$0.0 6.5
Reflection 0.0$\pm$0.0 6.5
Planning 0.0$\pm$0.0 6.5
Ref-Plan 0.0$\pm$0.0 6.5
GPT-4o Zero-shot 12.3$\pm$4.6 4
Reflection 26.2$\pm$19.4 1
Planning 23.2$\pm$14.2 3
Ref-Plan 23.6$\pm$22.1 2
@@
GPT-4o Text 23.6$\pm$22.1 3.5
Both 23.6$\pm$22.1 3.5
Gemini Text 51.9$\pm$31.9 1
Both 26.2$\pm$19.4 2
Claude Text 15.0$\pm$0.0 5
Both 9.7$\pm$4.6 6
@@
Environment. Baba Is You is a puzzle game in which players must discover and understand every rule and mechanic on their own, apart from the basic movement keys (‘left’, ‘right’, ‘up’, and ‘down’) [64]. The game’s defining feature is that the text tiles forming the rules can be pushed around, allowing the player to rewrite those rules on the fly. Every valid rule sentence must contain a verb (e.g., ‘Is’ and ‘Has’), and text tiles with a colored background (e.g., ‘You’, ‘Push’ and ‘Win’) cannot serve as subjects. A level is cleared when the object designated by ‘You’ touches the object designated by ‘Win’. For instance, with the rules ‘Baba Is You’ and ‘Flag Is Win’, the player wins as soon as Baba touches the flag. For implementation, we instrument the Steam edition with a lightweight Lua mod that, after every player move, dumps the full internal game state, i.e., the coordinates of every object, to a JSON file. The Lua modding script uses mod hook functions provided by the game developer, which are triggered at specific points in the game’s code. The agent outputs are delivered to the game via simulated key presses using the pyautogui library. We use Level 1 - Where do I go?, shown in Figure 22, as our default evaluation setting (for Table 3).
@@
$\text{Score}=\begin{cases}100,&\text{if level is cleared}\\ 40,&\text{if `Wall Is Stop' is broken and `Win' rule is created}\\ 20,&\text{if `Wall Is Stop' is broken only}\end{cases}$
@@
Llama-3.2-1B 6.7$\pm$11.5 12
Llama-3.2-3B 20.0$\pm$0.0 6.5
Qwen-2.5-3B 13.3$\pm$11.5 10.5
Qwen-2.5-7B 20.0$\pm$0.0 6.5
Minitron-4B 20.0$\pm$0.0 6.5
Minitron-8B 20.0$\pm$0.0 6.5
GPT-4o-mini 13.3$\pm$11.5 10.5
GPT-4o 20.0$\pm$0.0 6.5
o3-mini 73.3$\pm$46.2 1.5
Gemini-2.5-pro 73.3$\pm$46.2 1.5
Claude-3.7 46.7$\pm$46.2 3
Deepseek-R1 20.0$\pm$0.0 6.5
@@
Llama-3B Zero-shot 20.0$\pm$0.0 4.5
Reflection 20.0$\pm$0.0 4.5
Planning 20.0$\pm$0.0 4.5
Ref-Plan 20.0$\pm$0.0 4.5
GPT-4o Zero-shot 20.0$\pm$0.0 4.5
Reflection 20.0$\pm$0.0 4.5
Planning 20.0$\pm$0.0 4.5
Ref-Plan 20.0$\pm$0.0 4.5
@@
GPT-4o Text 20.0$\pm$0.0 6
Image 6.7$\pm$13.7 9
Both 20.0$\pm$0.0 6
Gemini Text 73.3$\pm$46.2 2
Image 20.0$\pm$0.0 6
Both 86.7$\pm$23.1 1
Claude Text 46.7$\pm$46.2 3
Image 20.0$\pm$0.0 6
Both 20.0$\pm$0.0 6
@@
Environment. 2048 [65] is a single-player sliding tile puzzle game played on a 4×4 grid. The objective is to combine numbered tiles by sliding them in one of four directions (i.e., up, down, left, or right) to create a tile with the value 2048. In this environment, the agent observes the current board state, represented as a 4×4 matrix of integers (each cell contains 0 for empty or a power of 2 for active tiles), and selects one of four discrete actions corresponding to directional moves. While the original game ends when no moves are available (i.e., the board is full and no adjacent tiles can be merged), we additionally terminate the episode if the agent performs five consecutive invalid moves (i.e., actions that result in no change to the board state). For implementation, we use an open-source, Pygame-based game environment. The logic module manages the board state, tile movements, merging, and win/loss conditions, while the interface leverages Pygame to render the board and handle user input. The implementation supports dynamic resizing and configurable parameters, and is designed to facilitate both human play and automated experiments.
@@
All reported results in Tables 4646 are computed as the mean and standard deviation over five independent runs, using a ‘zero-shot’ agent for each model configuration. To establish a baseline for comparison, we additionally included a random agent that selects one of the four possible actions (up, down, left, right) uniformly at random. This agent was evaluated over 50 episodes, and its performance is summarized in Tables.
@@
Random Agent 5.5$\pm$2.3 6
Llama-3.2-1B 0.0$\pm$0.1 15
Llama-3.2-3B 0.3$\pm$0.2 12
Qwen-2.5-3B 0.1$\pm$0.1 13
Qwen-2.5-7B 0.6$\pm$0.4 11
Minitron-4B 0.1$\pm$0.0 14
Minitron-8B 0.7$\pm$0.7 10
GPT-4o-mini 1.1$\pm$1.0 9
GPT-4o 5.6$\pm$1.5 5
o3-mini 25.3$\pm$7.3 2
o4-mini 15.7$\pm$6.7 3
o3 34.9$\pm$23.4 1
Gemini-2.5-pro 5.1$\pm$2.5 8
Claude-3.7 5.3$\pm$2.7 7
Deepseek-R1 11.5$\pm$3.4 4
@@
Random Agent - 5.5$\pm$2.3 6
Llama-3B Zero-shot 0.3$\pm$0.2 8
Reflection 0.0$\pm$0.0 11
Planning 0.1$\pm$0.1 10
Ref-Plan 0.1$\pm$0.2 9
GPT-4o Zero-shot 5.6$\pm$1.5 5
Reflection 3.5$\pm$2.9 7
Planning 6.0$\pm$5.5 4
Ref-Plan 7.0$\pm$5.7 3
o3-mini Zero-shot 25.3$\pm$7.3 1
Ref-Plan 17.6$\pm$9.5 2
@@
Random Agent - 5.5$\pm$2.3 5
GPT-4o Text 5.6$\pm$1.5 3
Image 1.8$\pm$1.1 10
Both 5.4$\pm$4.5 6
Gemini Text 5.1$\pm$2.5 8
Image 5.5$\pm$2.4 4
Both 3.1$\pm$2.6 9
Claude Text 5.3$\pm$2.7 7
Image 8.4$\pm$4.0 1
Both 6.7$\pm$0.9 2
@@
To increase the diversity of our training dataset and mitigate overfitting risks, we applied a data augmentation strategy focused on the system prompt. The original dataset was constructed by rolling out the same game prompt multiple times in the environment to collect assistant responses. The system prompt remained static, and the user prompt exhibited some variation due to changes in game state, though these were constrained to a fixed format where only specific values changed.
@@
Illustrative Example. Consider a multi-agent scenario in Street Fighter III. Agent 1 selects a ‘Super Attack’ that spans 7 frames, while Agent 2 chooses a ‘Move Away’ action that lasts 4 frames. The game orchestrator initiates both actions simultaneously. After 4 frames, Player 2’s action concludes, triggering the agent to observe the updated state and select a new action, e.g., a ‘Medium Punch’ lasting 2 frames. The orchestrator then integrates this new action into the ongoing simulation, even as Agent 1’s "Super Attack" continues. This frame-by-frame orchestration proceeds iteratively. Each agent re-enters the decision-making process immediately upon completing its current action, independent of the progress or status of other agents.
@@
Training Configuration. We fine-tune two models: Llama-3.2-1B and Llama-3.2-3B. Training is conducted using 4 NVIDIA A100 GPUs with 80GB of memory each. We use a learning rate of 1e-6, a per-device batch size of 4 with gradient accumulation steps set to 4, resulting in an effective batch size of 64. The models were trained for 1 epoch with 100 warm-up steps.
@@
Super Mario. For Supermario, we used Stage 3-1 as an unseen scenario, while we used Stage 1-1 for both the main evaluation and fine-tuning dataset. Unlike Stage 1-1, Stage 3-1 features different obstacle placements and introduces flying monster Koopas, making it a more challenging task. The game mechanics and control scheme remained unchanged.
@@
Appendix Q Limitations and Future Work
Cost Consideration. Among all games in Orak, six games, i.e., Ace Attorney, Her Story, Darkest Dungeon, Stardew Valley, Slay the Spire, and Baba is You, require a one-time purchase, typically priced ranging from \$$\$9.99$ to \$$\$24.99$. While this represents a non-negligible upfront cost, it is relatively minor compared to the recurring cost associated with proprietary LLM API calls. From a cost-efficiency perspective, the benchmark remains accessible and practical for sustained research.

### v2

@@ I.5 Result for Minecraft
Llama-3.2-1B $\pm0.0$ 9.5
Llama-3.2-3B $\pm0.0$ 9.5
Qwen-2.5-3B $\pm0.0$ 9.5
Qwen-2.5-7B $\pm0.0$ 9.5
Minitron-4B $\pm0.0$ 9.5
Minitron-8B $\pm0.0$ 9.5
[…]
@@
Llama-3B Zero-shot $\pm0.0$ 6
Reflection $\pm0.0$ 6
Planning $\pm0.0$ 6
Ref-Plan $\pm0.0$ 6
GPT-4o Zero-shot $\pm0.0$ 6
[…]
@@
Environment. Stardew Valley (ConcernedApe, 2016) is a life simulation and farming role-playing game. The player can engage in a variety of daily activities such as farming, fishing, mining, foraging, and socializing with villagers.
@@
We run the game on the Steam platform and use the modding tool SMAPI (Pathoschild, 2025) to extract in-game states and implement custom actions. To send keyboard and mouse inputs, we use the pyautogui library on macOS and the AutoHotkey (AHK) library on Windows, following prior work (Tan et al., 2024). Since many in-game actions such as planting or watering crops consist of multiple low-level actions (e.g., move up/down/right/left, switch tool, use tool), we define a set of high-level actions to abstract these into semantically meaningful units. Each high-level action is mapped to a predefined sequence of keyboard inputs and implemented via SMAPI scripts with custom keyboard bindings. We provide the save point used for our experiment to ensure reproducibility.
@@
The objective in the Stardew Valley task is to maximize the amount of money earned during the first 13 in-game days, starting from Day 1. This period is a natural milestone in the game, as the Egg Festival takes place on Spring 13, where players can purchase high-reward crops such as Strawberry seeds. We evaluate performance based on the net profit earned by the end of Spring 13, calculated as the difference between the final gold amount and the initial amount of 200 gold. We normalize the score using a human expert baseline of 1156200=9561156-200=956 gold, which we assign a normalized score of 100. Formally, the normalized score is defined as
$\text{Score}=(x_{\text{final}}-x_{\text{start}})/(x_{\text{oracle}}-x_{\text{start}})=(x_{\text{final}}-x_{\text{start}})/956,$
where $x_{\text{final}}$ is the agent’s final gold amount, $x_{\text{start}}=200$ is the starting gold, and $x_{\text{oracle}}=1156$ is the human expert score.
@@
Llama-3.2-1B $\pm0.0$ 10.5
Llama-3.2-3B $\pm0.0$ 10.5
Qwen-2.5-3B $\pm0.0$ 10.5
Qwen-2.5-7B $\pm0.0$ 10.5
Minitron-4B $\pm0.0$ 10.5
Minitron-8B $\pm0.0$ 10.5
[…]
@@
Llama-3B Zero-shot $\pm0.0$ 6.5
Reflection $\pm0.0$ 6.5
Planning $\pm0.0$ 6.5
Ref-Plan $\pm0.0$ 6.5
[…]
@@
GPT-4o Text $\pm4.8 1$
Image $\pm0.0$ 8.5
[…]
Both $0±\pm6.$0 2
[…]
Image $\pm0.0$ 8.5
[…]
@@
Environment. StarCraft II is a real-time strategy game where players gather resources, construct buildings, train units, and command armies to defeat opponents. The environment features a partially observable map, requiring the agent to explore and gather information about the opponent’s actions. For implementation, we adopt the BurnySc2/python-sc2 environment (BurnySc2, 2017), a Python interface widely used in the reinforcement learning (RL) community. Note that the library supports the raw scripted interface without a graphics-based interface. The environment supports various official maps and game modes. For our default evaluation (as in Table 3), we use the ‘Ancient Cistern LE’ map with the agent playing as Protoss against the built-in AI bot (Zerg, Hard difficulty; employing a timing build order strategy). The environment allows testing across different races, maps, and difficulty settings. For instance, the ‘Babylon LE’ is used to assess intra-game generalization in Table 7.
@@
$\text{Score}=\frac{\text{Number of Wins}}{\text{Total Matches Played}}\times 100$
@@
Using these difference-based scores, we then compute Elo ratings for all agents following the Bradley-Terry model (Bradley & Terry, 1952), following the approach used in StreetFighter III multi-agent evaluation in Section C.3.
@@
Llama-3.2-1B $\pm0.0$ 9.5
Llama-3.2-3B $\pm0.0$ 9.5
Qwen-2.5-3B $\pm0.0$ 9.5
Qwen-2.5-7B $\pm0.0$ 9.5
Minitron-4B $\pm0.0$ 9.5
Minitron-8B $\pm0.0$ 9.5
[…]
Claude-3.7 $0±\pm5$7.7 4.5
Deepseek-R1 $0±\pm5$7.7 4.5
@@
Llama-3B Zero-shot $\pm0.0$ 6
Reflection $\pm0.0$ 6
Planning $\pm0.0$ 6
Ref-Plan $\pm0.0$ 6
GPT-4o Zero-shot $0±\pm5$7.7 2.5
Reflection $\pm0.0$ 6
Planning $0±\pm5$7.7 2.5
[…]
@@
[…]
Both $0±\pm5$7.7 5
[…]
Claude Text $0±\pm5$7.7 5
Both $0±\pm5$7.7 5
@@
The game runs on the Steam platform, and we use modding tools to enable communication between the game and an external agent. In particular, we use BaseMod (Bug Kiooeht, 2018a) and ModTheSpire (Bug Kiooeht, 2018b), along with a modified version of CommunicationMod (Forgotten Arbiter, 2019), which enables state extraction and action input via standard input/output streams. Our modified version also extracts detailed in-game information, including full card and relic descriptions.
@@
$\text{Score}=\text{(\# of Cleared Floors)}+\frac{50}{3}\times\text{(\# of Bosses Defeated)}$ (1)
@@
Llama-3.2-1B $\pm0.0$ 10
Llama-3.2-3B $\pm0.0$ 10
Qwen-2.5-3B $\pm0.0$ 10
[…]
Minitron-4B $\pm0.0$ 10
Minitron-8B $\pm0.0$ 10
[…]
@@
Llama-3B Zero-shot $\pm0.0$ 6.5
Reflection $\pm0.0$ 6.5
Planning $\pm0.0$ 6.5
Ref-Plan $\pm0.0$ 6.5
[…]
@@
[…]
@@
Environment. Baba Is You is a puzzle game in which players must discover and understand every rule and mechanic on their own, apart from the basic movement keys (‘left’, ‘right’, ‘up’, and ‘down’) (Hempuli, 2019). The game’s defining feature is that the text tiles forming the rules can be pushed around, allowing the player to rewrite those rules on the fly. Every valid rule sentence must contain a verb (e.g., ‘Is’ and ‘Has’), and text tiles with a colored background (e.g., ‘You’, ‘Push’ and ‘Win’) cannot serve as subjects. A level is cleared when the object designated by ‘You’ touches the object designated by ‘Win’. For instance, with the rules ‘Baba Is You’ and ‘Flag Is Win’, the player wins as soon as Baba touches the flag. For implementation, we instrument the Steam edition with a lightweight Lua mod that, after every player move, dumps the full internal game state, i.e., the coordinates of every object, to a JSON file. The Lua modding script uses mod hook functions provided by the game developer, which are triggered at specific points in the game’s code. The agent outputs are delivered to the game via simulated key presses using the pyautogui library. We use Level 1 - Where do I go?, shown in Figure 22, as our default evaluation setting (for Table 3).
@@
[…]
@@
[…]
Qwen-2.5-3B $3±\pm1$1.5 10.5
[…]
GPT-4o-mini $3±\pm1$1.5 10.5
[…]
@@
[…]
@@
[…]
@@
Environment. 2048 (Cirulli, 2014) is a single-player sliding tile puzzle game played on a 4×4 grid. The objective is to combine numbered tiles by sliding them in one of four directions (i.e., up, down, left, or right) to create a tile with the value 2048. In this environment, the agent observes the current board state, represented as a 4×4 matrix of integers (each cell contains 0 for empty or a power of 2 for active tiles), and selects one of four discrete actions corresponding to directional moves. While the original game ends when no moves are available (i.e., the board is full and no adjacent tiles can be merged), we additionally terminate the episode if the agent performs five consecutive invalid moves (i.e., actions that result in no change to the board state). For implementation, we use an open-source, Pygame-based game environment. The logic module manages the board state, tile movements, merging, and win/loss conditions, while the interface leverages Pygame to render the board and handle user input. The implementation supports dynamic resizing and configurable parameters, and is designed to facilitate both human play and automated experiments.
@@
All reported results in Tables $45$ are computed as the mean and standard deviation over five independent runs, using a ‘zero-shot’ agent for each model configuration. To establish a baseline for comparison, we additionally included a random agent that selects one of the four possible actions (up, down, left, right) uniformly at random. This agent was evaluated over 50 episodes, and its performance is summarized in Tables.
@@
[…]
Qwen-2.5-3B $\pm0.1$ 13
[…]
Minitron-8B $\pm0.7$ 10
[…]
@@
[…]
Reflection $\pm0.0$ 11
Planning $\pm0.1$ 10
[…]
@@
[…]
Both $\pm4.5$ 6
[…]
@@
Prompting Details. To increase the diversity of our training dataset and avoid overfitting risks, we applied a data augmentation strategy focused on the system prompt. The original dataset was constructed by rolling out the same game prompt multiple times in the environment to collect assistant responses. The system prompt remains static, and the user prompt exhibits variations due to changes in game state, though these were constrained to a fixed format where only specific values changed.
@@
#\# of Augs. SuperMario 2048
0 24.2 1.4
3 26.8 3.5
10 26.7 2.8
20 25.9 2.4
Table 46: Performance across different numbers of augmentations.
Effect of The Number of Augmentations. Table 46 shows the fine-tuned model performance across different numbers of augmentations. We use the same fine-tuning setup as the intra-game generalization in Table 7. Overall, increasing the number of augmentations improves the performance of fine-tuned models compared to the one with no augmentation. Interestingly, the small number of augmentations (3 or 10) is quite effective to ensure linguistic diversity in fine-tuning, and the effectiveness diminishes as the number of augmentations grows.
@@
Illustrative Example. Consider a multi-agent scenario in Street Fighter III. Agent 1 selects a ‘Super Attack’ that spans 7 frames, while Agent 2 chooses a ‘Move Away’ action that lasts 4 frames. The game orchestrator initiates both actions simultaneously. After 4 frames, Player 2’s action concludes, triggering the agent to observe the updated state and select a new action, e.g., a ‘Medium Punch’ lasting 2 frames. The orchestrator then integrates this new action into the ongoing simulation, even as Agent 1’s ”Super Attack” continues. This frame-by-frame orchestration proceeds iteratively. Each agent re-enters the decision-making process immediately upon completing its current action, independent of the progress or status of other agents.
@@
Task LLaMA-3.2-1B(h) LLaMA-3.2-3B(h)
Intra-game/Non-game 1.7 4.2
OOD-game 1.7 4.3
Table 47: Elapsed GPU time for fine-tuning.
Training Configuration. We fine-tune two models: Llama-3.2-1B and Llama-3.2-3B. Training is conducted using 4 NVIDIA A100 GPUs with 80GB of memory each. We use a learning rate of 1e-6, a per-device batch size of 4 with gradient accumulation steps set to 4, resulting in an effective batch size of 64. The models were trained for 1 epoch with 100 warm-up steps. Table 47 summarizes the GPU time taken for fine-tuning.
@@
Appendix Q Analysis of The Gap between In-game and OOD-games in Fine-tuning
Genre Action Adventure RPG Simulation Strategy Puzzle
Games SuperMario AceAttorney DarkestD MineCraft SlaySpire 2048
Llama-3.2-1B (Pretrain) 18.7 1.3 0.0 0.0 0.0 0.1
Llama-3.2-1B (Fine-tune) 23.5 5.5 0.0 0.0 0.0 0.3
Llama-3.2-3B (Pretrain) 31.8 4.6 47.5 0.0 0.0 0.1
Llama-3.2-3B (Fine-tune) 24.9 9.1 40.1 0.0 0.0 3.1
Avg. Improvement (%) 2.0 210.5 -7.8 0.0 0.0 1600.0
Table 48: Performance of LLaMA models with varying in-ood game gap.
Setup. For studying the in-OOD gap in generalization, we fine-tune Llama-3.2 only on Baba Is You. Other hyperparameter configurations are exactly the same as in Appendix P.
Results. As shown in Table 48, fine-tuning yields the most improvements for OOD games in the same genre, but can also provide benefits across different genres in some cases. For example, 2048, another puzzle game, showed the largest improvement after fine-tuning. Interestingly, Ace Attorney, though a different genre, also benefited from fine-tuning, likely due to shared requirements in logical reasoning (see qualitative examples below). In contrast, for games like Darkest Dungeon and Super Mario, which involve different gameplay dynamics and action spaces, the effect was mixed. Note that, as shown in Table 7, unlike the current experiment (fine-tuned on Baba Is You only), LLMs fine‑tuned on 10 games (without Super Mario) showed consistent improvements on Super Mario (OOD game). This implies our fine‑tuning set, constructed from a diverse set of games across genres, contains knowledge broadly beneficial for gameplay/decision-making and supports cross‑genre generalizability.
Appendix R Limitations and Future Work
Cost Consideration. Among all games in Orak, six games, i.e., Ace Attorney, Her Story, Darkest Dungeon, Stardew Valley, Slay the Spire, and Baba is You, require a one-time purchase, typically priced ranging from $\mathdollar 9.99$ to $\mathdollar 24.99$. While this represents a non-negligible upfront cost, it is relatively minor compared to the recurring cost associated with proprietary LLM API calls. From a cost-efficiency perspective, the benchmark remains accessible and practical for sustained research.
License Issue. We have made considerable efforts to comply with licensing requirements in designing our benchmark. (1) Users must purchase the commercial games themselves for evaluation. (2) We do not distribute any commercial game executables. (3) We do not modify any game assets during gameplay. Moreover, we explicitly state the following legal compliance guidelines “This project is strictly for research purposes. Researchers are required to use our framework solely as a benchmark for evaluating LLMs, and any models derived from our benchmark are strictly prohibited from commercial use”.